The height of a triangle is 2 more than the base. The area is 40 square feet. Find the base.
Answer by mananth(16946)

base = x
height = (x+2)
area =40
Area of triangle = 1/2 * base * height
1/2 * x*(x+2)=40
x(x+2)=40*2
x^2+2x=80
x^2+2x-80=0
x^2+10x-8x-80=0
x(x+10)-8(x+10)=0
(x+10)(x-8)=0
x=-10 OR 8
ignore negative
base = 8 feet
height = 10 feet